OpenStudy (anonymous):

How can I solve this with elimination? 4x=2y-10 x+2y=5

jimthompson5910 (jim_thompson5910):

You need to get 4x=2y-10 into standard form first. Standard form is Ax + By = C To do this, you subtract 2y from both sides. 4x=2y-10 4x-2y = -10

jimthompson5910 (jim_thompson5910):

We now have this system 4x-2y=-10 x+2y=5 Adding the equations gives us 4x-2y=-10 x+2y=5 ---------- 5x+0y = -5 which is really 5x = -5. Solve for x to get x = ????